# Tracing the impacts of Mount Pinatubo eruption on global climate using spatially-varying changepoint detection

### v1

@@ 4.2 Results
Our method consistently returns a much lower FPR than the 1D method in all cases. When a changepoint is present, the FPR of our method decreases as the signal increases, while the FPR of the 1D method stays almost constant regardless of the signal strength. FNR of the two methods is comparable when the signal is strong, but our method largely outperforms 1D under weaker signal. FNR is not defined when there is no changepoint. Both the 1D and our method achieve smaller RMSE when signal is stronger, but our method outperforms the 1D method across all signal strengths with significantly smaller RMSE. The advantage of our method in terms of RMSE is more pronounced when the signal is weaker. Since the 1D method does not provide confidence intervals, we are unable to compare the two methods in terms of their uncertainty quantification. However, the empirical coverage of our 95% credible interval seems to be close to the nominal level for all tested cases, though their median coverage seems slightly above the nominal level. Although there is no simple interpretation for the empirical coverage of Bayesian credible intervals as there is for confidence intervals, our results still suggest a positive view of our credible intervals. The lengths of the credible interval appear to be informative given that the maximum interval length is $60$. Unsurprisingly, the length of the credible interval gets wider as signal becomes weaker.
Figure 5 shows the results for both variance shift only and mean and variance shift combined. We set $\gamma_{0}$ to be 0 and 2 associated with zero and positive mean shift, and set $\sigma^{2}_{\gamma}$ to be 3 and 5, representing two different levels of variance shift. Again, our method significantly outperforms the 1D method in both the detection and estimation of changepoints. The 1D method appears to be very conservative in detecting variance only shift, resulting in a low FPR but an extremely high FNR. This implies that the 1D method is more sensitive to mean changes than variance changes. However, when testing for a combined mean and variance shift, the 1D method detects changepoints for all locations and tends to identify a changepoint at $t=M-1$ when there is actually no changepoint, resulting in a low FNR but an extremely high FPR. In contrast, our method consistently performs well in detecting changepoints both in variance only change and in mean and variance change. The RMSE, empirical coverage, and length of credible intervals show a similar pattern as in Figure 4. Once again, both the RMSE and the length of credible intervals decrease as the signal increases. The results for mean and variance shift seem to be very similar to the corresponding mean shift only results with the mean shift set to 2. This suggests that the changepoint detection and estimation by our model are driven mainly by the mean shift rather than the variance shift in the presence of both types of changes.
Overall, our method outperforms the 1D method by taking advantage of the spatial correlation in the changepoint process for spatially indexed time series. The importance of borrowing strength from neighboring locations becomes more pronounced and necessary, especially when the change signal is weaker.
@@
Fig. 4: Boxplots of RMSE, FPR, FNR, the empirical coverage probability and the length of 95% credible intervals under different mean shift signal strength. “ST” is our Spatial-Temporal model and “1D” is the univariate method from R package changepoint.
@@
Fig. 5: Boxplots of RMSE, FPR, FNR, the empirical coverage probability and the length of 95% credible intervals under four settings comprised of two $\gamma_{0}$ values representing mean change and two $\sigma^{2}_{\gamma}$ values representing variance change. $\gamma_{0}=0$ indicates no mean change. “ST” is our Spatial-Temporal model and “1D” is the univariate method from R package changepoint.
@@ 5 Impact of Mt Pinatubo Volcanic Eruption
We apply our proposed method to AOD and temperature data to detect changes that are possibly caused by the stratospheric aerosol injection from the Mt. Pinatubo eruption in June 1991.
5.1 Stratospheric aerosol data
Following the suggestions in Section 3.2, we applied Model (1) on the aerosol data with both $\sigma^{2}_{1}<\sigma^{2}_{2}$ and $\sigma^{2}_{1}>\sigma^{2}_{2}$ since we have no prior knowledge about whether the variance increases or decreases after changepoint. In addition, we also test the case where $\sigma^{2}_{1}=\sigma^{2}_{2},$ which means that the change is soley due to a mean shift. We set ${\bf s}_{0}$ to be at ($123.4375^{\circ}$, $19^{\circ}$), which is an average of the AOD on the grid surface bounded by $(120^{\circ}$, $127.5^{\circ}$) in longitude and $(15^{\circ}$, $22.5^{\circ}$) in latitude. Since this grid contains the location of Mt. Pinatubo at ($120^{\circ}$, $15^{\circ}$), we anticipate the change to approximately spread from this location under our model assumption. Therefore, ${\bf s}_{0}$ is expected to have the earliest changepoint among all grid points, satisfying the definition ${\bf s}_{0}=\arg\min_{{\bf s}}\tau({\bf s})$. We used a discrete uniform prior for $\tau_{0}$ as well as weakly informative priors for all other parameters. Table 1 compares the DIC value and the posterior mean of key parameter estimates for the three forms of variance shift models. We observe that the equal variance model achieves the lowest DIC value. Consistent with this finding, the estimated variance shift parameter $\sigma^{2}_{\gamma}$ for the variance increase/decrease assumptions is very small compared to $\sigma^{2}_{1}$/$\sigma^{2}_{2}$. Thus, we proceed our inference based on the assumption $\sigma_{1}^{2}=\sigma_{2}^{2}$, that is, the change is solely due to a mean shift.
@@ 5.1 Stratospheric aerosol data
Table 1: DIC values and posterior summary of key parameters under different model assumptions for AOD data. The credible intervals in the last column are derived under $\sigma^{2}_{1}=\sigma^{2}_{2}$.
@@
We report the 95% credible intervals of key parameters under the assumption of $\sigma_{1}^{2}=\sigma_{2}^{2}$ in Table 1. The significantly positive estimate of $\gamma_{0F}$ indicates an elevated aerosol level immdediately after the volcanic eruption, and the negative estimate for $\gamma_{1F}$ indicates the subsequent restoration of aerosol level post-eruption, consistent with the anticipated trend. The diffusion parameter for longitude, $\beta_{1}$, is not significantly different from 0. In contrast, there is strong evidence to show that the diffusion parameter for latitude, $\beta_{2}$, is substantially greater than 0. This observation aligns with the spatial pattern of changepoints predominantly driven by latitude, as depicted in Figure 2(a) and 6.
@@
We determine the estimated changepoints by taking the posterior mode of the floor function of $\min(M,\tau(\mathbf{s}))$ at each location. Figure 6(a) shows the heatmap of estimated changepoints. Our model detects a changepoint at all locations, with estimated values ranging from Jun 1991 to Sep 1991. The estimated changepoints show a pattern driven more by latitude than longitude, with the earliest changepoints predominantly occurring along latitudes $3.5^{\circ}$ through $34^{\circ}$. This pattern aligns with our exploratory data analysis and is consistent with existing literature, which reports that the Pinatubo aerosol layer circled the Earth in 21 days and had spread to latitudes around $30^{\circ}N$ and $10^{\circ}S$ in the same period (Self et al., 1996). McCormick and Veiga (1992); Stowe et al. (1992) also found the Pinatubo aerosols straddled the equator.
Figure 6(b) displays the estimated mean shift parameter, $\gamma_{0}({\bf s})$, which ranges from $+0.25$ to $+1.5$. The spatial pattern of the estimated mean change matches with that shown in Figure 2(b), indicating the successful capture of the spatial variability in mean change by our method.
@@
The time series of aerosol after preprocessing and the estimated $\mu_{1}$ and $\mu_{2}(\mathbf{s},t)$ from our model are shown in Figure 7. Our mean estimates exhibit the same pattern as the preprocessed data, where the time series close to equator jump first and tend to have a larger peak compared to those around the latitudes further north or south. These results, together with Figure 6, demonstrate that our model effectively captures the diffusion of the impact of the volcanic eruption on the AOD data. Specifically, the estimated changepoints for locations near the latitude of Mt. Pinatubo ($15^{\circ}$) coincide with the month of the eruption, while the changepoints for locations further north or south occur several months later. We also plot the residuals for the AOD series by location in the Supplement, which verifies that $\sigma_{1}^{2}=\sigma_{2}^{2}$ is a reasonable assumption.
@@
Fig. 6: (a) Heatmap of detected changepoints. Darker color indicates earlier change. (b) Heatmap of estimated change amount. Darker color indicates larger magnitude of change. The red diamond marks the location of Mt. Pinatubo.
@@
Fig. 7: Time series of (a) aerosol data after preprocessing and (b) posterior mean of $\mu_{1}$ and $\mu_{2}(\mathbf{s},t)$, zoomed in to years $1991-1993$ The color represents the latitude of their location. The blue, yellow and red vertical lines mark the average value of the changepoints in latitude bands $(60^{\circ}S-20^{\circ}S),(20^{\circ}S-20^{\circ}N),$ and $(20^{\circ}N-60^{\circ}N)$.
5.2 Latitudinal surface temperature data
@@
Unlike the clear signal of volcanic impact on stratospheric AOD data, the impact on surface temperature is visually indiscernible, as observed in Figure 1. Indeed, the impact was too subtle to be detected when we applied our method to the gridded temperature data as was done for AOD data. Due to the zonal feature of the stratospheric aerosols spreading, the impact of volcanic eruption on surface temperature is more noticeable in latitude bands, as mentioned in Section 2. Consequently, we use zonal mean temperatures to trace the impact of eruption.The mechanism by which aerosols ejected into the lower stratosphere affect surface temperature is a complex process. Unlike AOD, we do not expect the volcanic impact on surface temperatures to diffuse from the event location Robock and Matson (1983). Thus, we negate the influence of the distance from $\mathbf{s_{0}}$ on $\tau_{0}({\bf s})$ by setting $\boldsymbol{\beta}=0$ in the mean of $\log(\boldsymbol{\Delta})$.
@@
Similar to aerosol data, we apply Model (1) with $\sigma^{2}_{1}<\sigma^{2}_{2}$, $\sigma^{2}_{1}>\sigma^{2}_{2}$, and $\sigma^{2}_{1}=\sigma^{2}_{2}$ on surface temperature and compare the results of the three model in Table 2. Again, the mean shift only model ($\sigma^{2}_{1}=\sigma^{2}_{2}$) achieves the lowest DIC, and the estimated variance change parameter $\sigma^{2}_{\gamma}$ for the two variance shift models is very small compared to $\sigma^{2}_{1}$/$\sigma^{2}_{2}$. We thus proceed our analysis under the assumption $\sigma^{2}_{1}=\sigma^{2}_{2}$. Table 2 reports the posterior mean and the 95% credible interval of key parameters. Unlike the aerosol data, the mean shift parameter $\gamma_{0F}$ for temperature is significantly negative, indicating the global cooling effect of eruption consistent with other literature. The slope parameter $\gamma_{1F}$ is also significantly negative, indicating continued cooling of the global temperature during the time period considered in our data.
@@
Table 2: DIC values and posterior summary of key parameters under different model assumptions for surface temperature data. The credible intervals in the last column are derived under $\sigma^{2}_{1}=\sigma^{2}_{2}$.
Figure 8 shows the posterior mean of $\mu_{1}$ and $\mu_{2}({\bf s},t)$ for the temperature series at each latitude, separated at the posterior mode of $\tau({\bf s}).$ Our method detects changepoints for all latitudes in the range $(56^{\circ}S-49^{\circ}N)$ except for latitudes $18.5^{\circ}S,26.5^{\circ}N,$ and $34.0^{\circ}N.$ No changepoints were detected in the southern $(86^{\circ}S-63.5^{\circ}S)$ and northern $(56.5^{\circ}N-86.5^{\circ}N)$ ends of the globe. The posterior probability of having no changepoint ranges from 0 to 0.0112 for detected latitudes and 0.7955 to 0.9986 for the undetected latitudes. The detected changepoints range from Sep 1991 - May 1992. See Table 3 in the Supplement for further details.
@@
The magnitude of mean shift shown in Figure 8 is not true to scale since we normalized the data to make the variance constant for all locations in the preprocessing step. Figure 9 shows the posterior distribution of the actual mean shift at the detected latitudes, after back-transforming to match the scale of the original data. While there were more changepoints detected in the southern hemisphere, the amount of change is smaller compared to the northern hemisphere, and the change in global temperature is mostly driven by the latitudes $41.5^{\circ}N$ and $49^{\circ}N$. The average mean change in the southern hemisphere, northern hemisphere, and the globe is $-0.136^{\circ}C,-0.205^{\circ}C,$ and $-0.191^{\circ}C,$ respectively. Our findings are consistent with the existing studies (Self et al., 1996), which report that the Mt. Pinatubo eruption led to a global cooling of $0.4^{\circ}C$ between years $1991$ and $1993$, driven mostly by a cooling of $0.5^{\circ}C$ in Northern Hemisphere. The estimated change amounts returned by out model are smaller in magnitude, which is expected since our method only measures the amount of mean shift at the time of the changepoint as opposed to the average change over two years. The negative $\gamma_{1F}$ estimate also indicates that the temperature continues to cool down after the changepoint.
@@
Fig. 8: Posterior mean of $\mu_{1}$ and $\mu_{2}({\bf s},t)$ for surface temperature, separated at the posterior mode of $\tau({\bf s})$ (red lines). The black lines are latitudinal mean temperature series after preprocessing. The posterior distribution of the changepoints are overlayed at the bottom of each plot. No changepoints were detected at $86.0^{\circ}$ $63.5^{\circ}$, $18.5^{\circ}$, $26.5^{\circ}$, $34.0^{\circ}$, and $56.5^{\circ}$ $86.5^{\circ}$.
@@
Fig. 9: Posterior distribution of $\gamma_{0}({\bf s})$ (a) by latitude and (b) averaged over hemisphere after backtransforming to original scale. The red line on (a) marks the equator.
@@
Our findings from the analysis of stratospheric AOD and surface temperature data are consistent with the previous studies in general. However, our work complements the previous studies by providing quantitative measures of the impact of the Pinatubo eruption on aerosol and climate, as well as specifying the spatial pattern of the impact.

### v2

@@ 4.2 Results
The FPR of both methods are comparable when a changepoint is present. Still, our method seems more stable when signal is strong. When there is no changepoint, our method returns a much lower FPR. The FNRs of the two methods are comparable when the signal is strong, but our method largely outperforms the 1D under a weaker signal. Note that FNR is not defined when there is no changepoint. Both the 1D and our method achieve smaller RMSE when the signal is stronger, but our method outperforms the 1D method across all signal strengths with significantly smaller RMSE. The advantage of our method in terms of RMSE is more pronounced when the signal is weaker. The empirical coverage of the $95\%$ credible intervals obtained from the 1D method appears to be above the nominal level for all tested cases. However, the mean empirical coverage of our method is centered around the nominal level. More importantly, our method results in a much narrower and thus more informative credible interval compared to the 1D method. Unsurprisingly, the length of the credible interval gets wider as the signal becomes weaker. Although there is no simple interpretation for the empirical coverage of Bayesian credible intervals as there is for confidence intervals, our results still suggest a positive view of our credible intervals. The lengths of the credible interval appear to be informative given that the maximum interval length is $60$. Unsurprisingly, the length of the credible interval gets wider as the signal becomes weaker.
@@
Fig. 4: Boxplots of RMSE, FPR, FNR, the empirical coverage probability and the length of 95% credible intervals under different mean shift signal strength. “ST” is our spatio-temporal model and “1D” is the univariate method. FNR is not defined under Mean Shift=0.
Figure 5 shows the results for both variance shift only and mean and variance shift combined. We show results for $\gamma_{0}\in\{0,2\}$ associated with zero and positive mean shift, and the variance difference parameter $\sigma^{2}_{\gamma}\in\{3,5\}$, representing two different levels of variance shift. The results for $\gamma_{0}\in\{1.5,3\}$ and $\sigma^{2}_{\gamma}\in\{3,5\}$ are provided in Section 2 of the Supplement (Shi-Jun et al., 2023).
In the case of variance only shift, our method significantly outperforms the 1D method, with even a wider margin, in all metrics except for FPR. The 1D method appears to be very conservative in detecting variance only shift, resulting in a lower FPR but an extremely high FNR. This implies that the 1D method is less sensitive to variance changes than mean changes.
For combined mean and variance shift, the results seem to be very similar to those for mean shift only. The FPR is comparable for both methods, while the FNR and RMSE are much lower for our method. In this combined case, the empirical coverage of credible intervals for the 1D method is closer to, though still larger than, the nominal level. Once again, the credible intervals obtained from our method are much shorter than the 1D method without sacrificing their empirical coverage.
For each simulation, we calculate the posterior estimate of the parameter by taking posterior mode for $\tau_{0}$ and posterior mean for all other parameters. To give an example of parameter estimation with our method, we summarize the parameter estimates for the simulations with mean shift of 1.5 and $\sigma^{2}_{\gamma}=0$ in Table 1. The second and third columns of the table report the mean estimates and the $95\%$ credible intervals across the 100 simulations. We observe that the true value is captured by the $95\%$ interval for all parameters in the model.
@@
Fig. 5: Boxplots of RMSE, FPR, FNR, the empirical coverage probability and the length of 95% credible intervals under four settings comprised of two $\gamma_{0}$ values representing mean change and two $\sigma^{2}_{\gamma}$ values representing variance change. $\gamma_{0}=0$ indicates no mean change. “ST” is our spatio-temporal model and “1D” is the univariate method.
Parameter True Value Mean of the estimate 95% CI of the mean
$\alpha_{0}$ 0 0.0014 (-0.0831, 0.0754)
$\gamma_{0F}$ 1.5 1.4616 (1.3260, 1.6342)
$\sigma^{2}_{1}$ 1 1.0309 (0.8844, 1.2461)
$\sigma^{2}_{U}$ 1 0.9335 (0.7634, 1.1445)
$\psi_{U}$ 2 1.8901 (1.5467, 2.3630)
$\phi_{U}$ 1.5 1.4489 (1.1393, 1.7735)
$\tau_{0}$ 18 16 (8, 20)
$\beta_{1}$ 1.5 1.4366 (0.9382, 2.3505)
$\beta_{2}$ 1 1.0408 (0.6369, 1.6172)
$\sigma^{2}_{\Delta}$ 1 0.8644 (0.6362, 1.3412)
$\psi_{\Delta}$ 0.5 0.6572 (0.3544, 1.0177)
Table 1: Summary of parameter estimates for a mean shift of 1.5 and variance shift of 0.
Overall, our method outperforms the 1D method by taking advantage of the spatial correlation in the changepoint process for spatially indexed time series. The importance of borrowing strength from neighboring locations becomes more pronounced and necessary when the change signal is weaker. Similar phenomena were observed by Wang et al. (2023).
@@ 5 Impact of Mt Pinatubo Volcanic Eruption
We apply our proposed method to AOD and surface temperature data to detect changes that are possibly caused by the aerosols injected into the stratosphere by the Mt. Pinatubo eruption in June 1991. All MCMC convergence in the data analyses have been verified by the Gelman-Rubin diagnostic using three parallel chains with different initial values for the parameters.
5.1 Aerosol Optical Depth
Following the suggestions in Section 3.2, we apply model (1) on the aerosol data with both $\sigma^{2}_{1}<\sigma^{2}_{2}$ and $\sigma^{2}_{1}>\sigma^{2}_{2}$ since we have no prior knowledge about whether the variance increases or decreases after changepoint. In addition, we also test the case of $\sigma^{2}_{1}=\sigma^{2}_{2},$ which means that the change is solely due to a mean shift.
We set ${\bf s}_{0}$ to be at ($123.4375^{\circ}$E, $19^{\circ}$N), the center of the grid cell that contains the location of Mt. Pinatubo ($120^{\circ}$E, $15^{\circ}$N). We anticipate ${\bf s}_{0}$ has the earliest changepoints among all grid points. We use a discrete uniform prior for $\tau_{0}$ and weakly informative priors for all other parameters. Table 2 compares the DIC value and the posterior mean of key parameter estimates for the three forms of variance shift models. We observe that the equal variance model achieves the lowest DIC value. Consistent with this finding, the estimated variance shift parameter $\sigma^{2}_{\gamma}$ for the variance increase or decrease assumptions is very small compared to $\sigma^{2}_{1}$ or $\sigma^{2}_{2}$. Thus, we proceed with our inference based on the assumption $\sigma_{1}^{2}=\sigma_{2}^{2}$, that is, the change is only due to a mean shift.
@@ 5.1 Aerosol Optical Depth
$\sigma^{2}_{\Delta}$ 0.9036 0.8288 0.8587 (0.4766,1.786)
$\psi_{U}$ 2.5679 2.8244 2.6175 (2.0745,2.9091)
$\phi_{U}$ 0.8336 0.8965 0.8229 (0.7161,0.9198)
$\psi_{\Delta}$ 0.8084 0.8489 0.8212 (0.3817,1.7390)
$\sigma^{2}_{\gamma_{0}}$ 1.5794 1.3702 1.7418 (1.0757,2.1393)
$\sigma^{2}_{\gamma_{1}}$ 0.0006 0.0007 0.0007 (0.0004,0.0011)
$\psi_{\gamma_{0}}$ 6.7958 7.5673 5.8663 (4.1403,10.7769)
$\psi_{\gamma_{1}}$ 6.5380 5.9436 5.8552 (4.3915,9.9309)
@@
Table 2: DIC values and posterior summary of parameters under different model assumptions for AOD data. The credible intervals in the last column are derived under $\sigma^{2}_{1}=\sigma^{2}_{2}$.
@@
We report the 95% credible intervals of key parameters under the assumption of $\sigma_{1}^{2}=\sigma_{2}^{2}$ in Table 2. The significantly positive estimate of $\gamma_{0F}$ indicates an elevated aerosol level immdediately after the volcanic eruption, and the negative estimate of $\gamma_{1F}$ indicates the subsequent restoration of aerosol level post-eruption, consistent with the anticipated trend. The diffusion parameter for longitude, $\beta_{1}$, is not significantly different from 0. In contrast, there is strong evidence to show that the diffusion parameter for latitude, $\beta_{2}$, is substantially greater than 0. This observation aligns with the spatial pattern of changepoints predominantly driven by latitude, as depicted in Figure 2(a).
@@
We determine the estimated changepoints by taking the posterior mode of the floor function of $\min\{M,\tau(\mathbf{s})\}$ at each location. Figure 6(a) shows the heatmap of estimated changepoints. Our model detects a changepoint at all locations, with estimated values ranging from Jun 1991 to Sep 1991. The estimated changepoints show a pattern driven more by latitude than longitude, with the earliest changepoints predominantly occurring along latitudes $3.5^{\circ}$S through $34^{\circ}$N. This pattern aligns with our exploratory data analysis and is consistent with existing literature, which reports that the Pinatubo aerosol layer circled the Earth in 21 days and had spread to latitudes around $30^{\circ}N$ and $10^{\circ}S$ in the same period (Self et al., 1996). McCormick and Veiga (1992) and Stowe et al. (1992) also found the Pinatubo aerosols straddled the equator. Figure 6(b) displays the estimated mean shift parameter, $\gamma_{0}({\bf s})$, which ranges from +0.25+0.25 to +1.5+1.5. The spatial pattern of the estimated mean change matches with that shown in Figure 2(b), indicating the successful capture of the spatial variability in mean change by our method.
@@
The time series of AOD after preprocessing and the estimated $\mu_{1}$ and $\mu_{2}(\mathbf{s},t)$ from our model are shown in Figure 7. Our mean estimates follow the same pattern as the preprocessed data, of which the time series close to the equator in the latitude bands $20^{\circ}S-20^{\circ}N$ jump first and tend to have a higher peak compared to those around the latitudes further north or south. These results, together with Figure 6, demonstrate that our model effectively captures the diffusion of the impact of the volcanic eruption on AOD. Specifically, the estimated changepoints for locations near the latitude of Mt. Pinatubo ($15^{\circ}$N) coincide with the month of the eruption, while the changepoints for locations further north or south occur several months later. We also plot the residuals for the AOD series by location in Section 2 of the Supplement (Shi-Jun et al., 2023), which verifies that $\sigma_{1}^{2}=\sigma_{2}^{2}$ is a reasonable assumption.
@@
Fig. 6: (a) Heatmap of detected changepoints for AOD. Darker color indicates earlier change. (b) Heatmap of estimated change amount. Darker color indicates larger magnitude of change. The white diamond marks the location of Mt. Pinatubo.
@@
Fig. 7: Time series of (a) AOD after preprocessing and (b) posterior mean of $\mu_{1}$ and $\mu_{2}(\mathbf{s},t)$, zoomed in to years 19911993.1991-1993. The dashed line marks the average value of the changepoints in latitude bands $(60^{\circ}S-20^{\circ}S),(20^{\circ}S-20^{\circ}N),$ and $(20^{\circ}N-60^{\circ}N)$.
5.2 Zonal surface temperature data
@@
Unlike the clear signal of volcanic impact on stratospheric AOD data, the impact on surface temperature is visually indiscernible, as observed in Figure 1. Indeed, the impact was too subtle to be detected when we applied our method to the gridded temperature data as was done for AOD data. Due to the strong latitudinal, or zonal, trend of AOD, the impact of volcanic eruption on surface temperature is not as local as the observed impact on AOD. Changes in surface temperature following the eruption are more noticeable when aggregated by zonal bands, as mentioned in Section 2. Consequently, we use zonal mean temperatures to trace the impact of the eruption. The mechanism behind how aerosols ejected into the lower stratosphere affect surface temperature is a complex process. Unlike AOD, we do not expect the volcanic impact on surface temperatures to diffuse from the event location (Robock and Matson, 1983). Thus, we negate the influence of the distance from $\mathbf{s_{0}}$ on $\tau_{0}({\bf s})$ by setting $\boldsymbol{\beta}=0$ in the mean of $\log(\boldsymbol{\Delta})$. For this data, ${\bf s}_{0}$ is interpreted as the location corresponding to the smallest $\tau({\bf s})$ estimate.
@@
Similar to the AOD data, we apply Model (1) with $\sigma^{2}_{1}<\sigma^{2}_{2}$, $\sigma^{2}_{1}>\sigma^{2}_{2}$, and $\sigma^{2}_{1}=\sigma^{2}_{2}$ to surface temperature and compare the results of the three models in Table 3. Again, the mean shift only model ($\sigma^{2}_{1}=\sigma^{2}_{2}$) achieves the lowest DIC, and the estimated variance change parameter $\sigma^{2}_{\gamma}$ for the two variance shift models is very small compared to $\sigma^{2}_{1}$ or $\sigma^{2}_{2}$. We thus proceed with our analysis under the assumption $\sigma^{2}_{1}=\sigma^{2}_{2}$. Table 3 reports the posterior mean and the 95% credible interval of key parameters. Unlike the AOD data, the mean shift parameter $\gamma_{0F}$ for temperature is significantly negative, indicating the global cooling effect of eruption consistent with other literature. The slope parameter $\gamma_{1F}$ is also significantly negative, indicating continued cooling of the global temperature after the changepoint during the time period considered in our data.
@@
$\sigma^{2}_{\Delta}$ 20.16 20.92 20.43 (8.33,56.65)
$\psi_{U}$ 4.9675 4.8348 4.8779 (4.2658,5.6670)
$\phi_{U}$ 1.2129 1.2063 1.1958 (1.0662,1.3851)
$\psi_{\Delta}$ 1.7805 2.0915 1.8195 (0.8843,3.0968)
$\sigma^{2}_{\gamma_{0}}$ 0.0390 0.03870 0.0312 (0.0145,0.1516)
$\sigma^{2}_{\gamma_{1}}$ 3.49e-5 3.63e-5 3.56e-5 (1.93e-6,9.88e-5)
$\psi_{\gamma_{0}}$ 21.46 20.63 23.38 (15.29,28.47)
$\psi_{\gamma_{1}}$ 27.48 25.75 25.51 (18.50,28.66)
@@
Table 3: DIC values and posterior summary of parameters under different model assumptions for surface temperature data. The credible intervals in the last column are derived under $\sigma^{2}_{1}=\sigma^{2}_{2}$.
Figure 8 shows the posterior mean of $\mu_{1}$ and $\mu_{2}({\bf s},t)$ for the temperature series at each latitude, separated at the posterior mode of $\tau({\bf s}).$ Our method detects changepoints for all latitudes in the range $(56^{\circ}S-49^{\circ}N)$ except for latitudes $18.5^{\circ}S,26.5^{\circ}N,$ and $34.0^{\circ}N.$ No changepoints were detected in the southern $(86^{\circ}S-63.5^{\circ}S)$ and northern $(56.5^{\circ}N-86.5^{\circ}N)$ ends of the globe. The posterior probability of having no changepoint ranges from 0 to 0.0112 for the detected latitudes and 0.7955 to 0.9986 for the undetected latitudes. The detected changepoints range from Sep 1991 - May 1992. See Table 1 in the Supplement (Shi-Jun et al., 2023) for further details. The lack of changepoint in the latitudes $18.5^{\circ}S,26.5^{\circ}N,34.0^{\circ}N$ may potentially be due to low signal-to-noise ratio in the subtropics (roughly 23 to 35 north and south) which are particulary synoptically active (McClain et al., 2004; Ryoo et al., 2008; Lensky et al., 2018). As shown in Figure 8, at all these three latitudes, we observe a noticeable drop in the temperature shortly after the Pinatubo eruption but then immediately followed by a rise, likely due to the volatile nature of subtropics. The few data points in the temperature drop make the changepoint detection very challenging as they simply behave like noise. In conclusion, we cannot assert that these three latitudes were not affected by the Mt. Pinatubo eruption, but the fingerprint of the eruption on those latitudes (if any) is brief and weak that can be easily masked by the natural fluctuation, based on the MERRA-2 data we used in our analyses.
@@
The magnitude of the mean shift shown in Figure 8 is not true to scale since we normalized the data to make the variance constant for all locations in the preprocessing step. Figure 9 shows the posterior distribution of the actual mean shift at the detected latitudes, after back-transforming to match the scale of the original data. While there were more changepoints detected in the southern hemisphere, the magnitude of change is smaller compared to the northern hemisphere, and the change in global temperature is mostly driven by the latitudes $41.5^{\circ}N$ and $49^{\circ}N$. The average mean change in the southern hemisphere, northern hemisphere, and the globe is $0.136^{\circ}C,-0.205^{\circ}C,$ and $0.191^{\circ}C,$ respectively. Our findings are consistent with the existing studies (Self et al., 1996), which report that the Mt. Pinatubo eruption led to a global cooling of $0.4^{\circ}C$ between years $1991$ and $1993$, driven mostly by a cooling of $0.5^{\circ}C$ in Northern Hemisphere. The estimated change amounts by our model are smaller in magnitude, which is expected since our method only measures the amount of mean shift at the time of the changepoint as opposed to the average change over two years. The negative $\gamma_{1F}$ estimate also indicates that the temperature continues to cool down after the changepoint.
@@
Fig. 8: Posterior mean of $\mu_{1}$ and $\mu_{2}({\bf s},t)$ for surface temperature, separated at the posterior mode of $\tau({\bf s})$ (thick black lines). The thin black lines are latitudinal mean temperature series after preprocessing. The posterior distribution of the changepoints are overlayed at the bottom of each plot. No changepoints were detected at $86.0^{\circ}$S - $63.5^{\circ}$S, $18.5^{\circ}$S, $26.5^{\circ}$N, $34.0^{\circ}$N, and $56.5^{\circ}$N - $86.5^{\circ}$N. The dashed vertical lines mark the month of the eruption (1991/06)
@@
Fig. 9: Posterior distribution of $\gamma_{0}({\bf s})$ (a) by latitude and (b) averaged over hemisphere after backtransforming to the original scale of surface temperature data. The dashed line in (a) marks the equator.
@@
Our findings from the analysis of stratospheric AOD and surface temperature data are consistent with the previous studies in general. However, our work complements the previous studies by providing quantitative regional measures of the impact of the Pinatubo eruption on aerosol and climate, as well as specifying the spatial pattern of the impact.